OpenStudy (anonymous):

simplify the equation: (-6i + 7i) + (6 - 7i)

OpenStudy (nikita2):

6-6i

OpenStudy (anonymous):

add real part with real part and imaginary with imaginary

OpenStudy (anonymous):

I made a mistake it's (-6 + 7i) + (6 - 7i)

OpenStudy (zzr0ck3r):

\(=-6+6+7i-7i=0+0=0\)

OpenStudy (anonymous):

so it's just 0 I understand I just didn't think that could be an answer thanks